# Few-Shot Anomaly Detection via Category-Agnostic Registration Learning

### v1

@@ VI-G Cross-dataset Generalization
The detailed results of anomaly detection/localization for each category can be found in Table X, offering insights into anomaly detection and localization outcomes under the PC estimator. It serves to validate the robustness of registration learning. This is evident from the fact that even when the distribution of the training data is altered, a significant improvement compared to the baseline is sustained. For instance, when evaluating on MVTec with K=2, CAReg demonstrates a performance enhancement of 4.4%/2.5% when trained on its original dataset (MVTec). Even when transitioning to a distinct distribution (MPDD) for training, the improvement remains substantial at 3.6%/1.6%. In essence, the results firmly establish the proposed method’s ability to generalize across datasets, showcasing how the influence of training with an entirely different dataset is minimal due to the inherent category-agnostic nature of the registration training approach.
@@ VI-H Visualization Analysis
Figure 6 provides qualitative analysis results to showcase the effectiveness of the category-agnostic feature registration training approach and its impact on anomaly detection and localization performance. With the PaDim normal distribution estimator, the anomaly localization produced by CAReg (column e) is observed to be closer to the ground truth (GT) (column j) compared to its corresponding baseline (column c). Similarly, the results produced by CAReg with the PatchCore normal distribution estimator (column i) are closer to GT than those produced by its corresponding baseline (column g). These visualizations illustrate the effectiveness of the training procedure via feature registration aggregated on multiple categories, resulting in improved anomaly localization.
In Figure 7, t-SNE [85] is used to visualize the features extracted on the MVTec test set, with each dot representing a normal image. It is evident that after the feature registration training, the features become more compact within each category. Simultaneously, features from different categories are better separated. The presence of more compact feature distributions is desirable, as it enhances the accuracy of the normal distribution estimate for each category. This improved separation of features contributes to better discrimination between normal and anomalous samples, leading to enhanced AD performance. Overall, the visualization results provide qualitative evidence of the efficacy of the category-agnostic feature registration training approach in enhancing anomaly detection and localization.
@@ VII Conclusion
This paper introduces a novel training method, called CAReg, for few-shot anomaly detection (FSAD), by learning generalizable registration techniques across different categories using only normal images for each category. This allows the model to accurately register a test image with its corresponding support (normal) images from unseen novel categories without the need for re-training or parameter fine-tuning. The experimental results demonstrate that CAReg outperforms state-of-the-art FSAD methods in both anomaly detection and anomaly localization tasks on standard benchmark datasets. Even when compared to anomaly detection methods trained with much larger volumes of data, CAReg remains competitive. It significantly enhances both the accuracy and efficiency of FSAD, showcasing its high potential for real-world anomaly detection applications.
Acknowledgments. This work was supported by the National Key Research and Development Program of China (No. 2020YFB1406801), 111 plan (No. BP0719010), STCSM (No. 18DZ2270700), State Key Laboratory of UHD Video and Audio Production and Presentation, and NUS Faculty Research Committee Grant (WBS: A-0009440-00-00).

### v2

@@ VI-G Cross-dataset Generalization
The detailed results of AD/localization for each category can be found in Table X, offering insights into AD and localization outcomes under the PC estimator. It serves to validate the robustness of registration learning. This is evident from the fact that even when the distribution of the training data is altered, a significant improvement compared to the baseline is sustained. For instance, when evaluating on MVTec with K=2, CAReg demonstrates a performance enhancement of 4.4%/2.5% when trained on its original dataset (MVTec). Even when transitioning to a distinct distribution (MPDD) for training, the improvement remains substantial at 3.6%/1.6%. In essence, the results firmly establish the proposed method’s ability to generalize across datasets, showcasing how the influence of training with an entirely different dataset is minimal due to the inherent category-agnostic nature of the registration training approach.
@@ VI-H Visualization Analysis
Fig. 6 provides qualitative analysis results to showcase the effectiveness of the category-agnostic feature registration training approach and its impact on AD and localization performance. With the PaDim normal distribution estimator, the anomaly localization produced by CAReg (column e) is observed to be closer to the ground truth (GT) (column j) compared to its corresponding baseline (column c). Similarly, the results produced by CAReg with the PatchCore normal distribution estimator (column i) are closer to GT than those produced by its corresponding baseline (column g). These visualizations illustrate the effectiveness of the training procedure via feature registration aggregated on multiple categories, resulting in improved anomaly localization.
In Fig. 7, t-SNE [85] is used to visualize the features extracted on the MVTec test set, with each dot representing a normal image. It is evident that after the feature registration training, the features become more compact within each category. Simultaneously, features from different categories are better separated. The presence of more compact feature distributions is desirable, as it enhances the accuracy of the normal distribution estimate for each category. This improved separation of features contributes to better discrimination between normal and anomalous samples, leading to enhanced AD performance. Overall, the visualization results provide qualitative evidence of the efficacy of the category-agnostic feature registration training approach in enhancing AD and localization.
@@ VII Conclusion
This paper introduces a novel training method, called CAReg, for FSAD, by learning generalizable registration techniques across different categories using only normal images for each category. This allows the model to accurately register a test image with its corresponding support (normal) images from unseen novel categories without the need for re-training or parameter fine-tuning. The experimental results demonstrate that CAReg outperforms state-of-the-art FSAD methods in both AD and anomaly localization tasks on standard benchmark datasets. Even when compared to AD methods trained with much larger volumes of data, CAReg remains competitive. It significantly enhances both the accuracy and efficiency of FSAD, showcasing its high potential for real-world AD applications.